OpenStudy (anonymous):

could you solve x^2x+1=75

OpenStudy (anonymous):

x^3 + 1 = 75

hero (hero):

Actually, I can't tell whether its: x^(2x) + 1 = 75 or x^2 * x + 1 = 75

OpenStudy (anonymous):

i keep saying this if you have question SIMILAR to this use http://www.wolframalpha.com/bing/?i=x%5e2x%2b1%3d75&expl=3

OpenStudy (mathmate):

If it is x^(2x)+1=75 (x^x)^2=(75-1) x^x = sqrt(74) x=2.7063 approx.